OpenStudy (anonymous):

If AD = 7 and AC = 28, calculate the length of line segment AB.

OpenStudy (anonymous):

If BD = 0.6 in. and AB = 0.8 in., calculate the length of line segment DA. .6^2 + .8^2 = DA^2 .36 + .64 = DA^2 1 = DA^2 (take square root) 1=DA